# Generative Human Geometry Distribution

### v2

@@ Abstract
Realistic human geometry generation is an important yet challenging task, requiring both the preservation of fine clothing details and the accurate modeling of clothing-pose interactions. Geometry distributions, which can model the geometry of a single human as a distribution, provide a promising representation for high-fidelity synthesis. However, applying geometry distributions for human generation requires learning a dataset-level distribution over numerous individual geometry distributions. To address the resulting challenges, we propose a novel 3D human generative framework that, for the first time, models the distribution of human geometry distributions. Our framework operates in two stages: first, generating the human geometry distribution, and second, synthesizing high-fidelity humans by sampling from this distribution. We validate our method on two tasks: pose-conditioned 3D human generation and single-view-based novel pose generation. Experimental results demonstrate that our approach achieves the best quantitative results in terms of realism and geometric fidelity, outperforming state-of-the-art generative methods.
Figure 1: (a) Our method synthesizes high-fidelity human geometry using a diffusion denoising model. Building on this model, we propose a versatile generative framework and perform two generative tasks. (b) Pose-conditioned random generation: given a human pose, we randomly synthesize realistic geometry with natural clothing details. (c) Novel pose generation: given a normal image, we synthesize high-quality geometry in novel poses. Our results showcase detailed and pose-aware clothing deformations. All visual results are rendered from point clouds.
@@ 1 Introduction
3D human geometry generation is a critical yet challenging task. The human body exhibits high-frequency clothing details, which are inherently difficult to synthesize. Therefore, the core of this task is the design of a representation that can capture both underlying structure and fine details, posing two key challenges: 1) encoding high-frequency details into a low-dimensional manifold without losing fidelity; 2) modeling the relationship between clothing wrinkles and poses to preserve realistic details.
@@
Existing methods represent the human body in different ways. NeRFs [23, 53, 38, 14, 45, 58] focus on the rendering results but neglect underlying geometry and are constrained by the rendering speed and resolution. Implicit functions [41], such as signed distance functions struggle to synthesize thin structures and tend to oversmooth results. Point clouds [55, 56] and volume-based representations trade-off between memory efficiency and quality. Recently, geometry distributions [52] model each geometry as a distribution, enabling “infinite” sampling from the distribution to construct high-fidelity geometry, offering a promising approach for detailed human representation. However, extending a single-geometry distribution, trained on a single shape, to an entire dataset is non-trivial and computationally inefficient, as it requires modeling the dataset as a distribution over numerous geometry distributions, as illustrated in Fig. 2, introducing hierarchical complexity.
Figure 2: The dataset distribution represents the distribution of geometries within the dataset, where each sample corresponds to a geometry. The geometry distribution illustrates the distribution of points within a single geometry. Most generative methods model the dataset distribution only, while our method models the dataset distribution of geometry distributions.
To this end, we propose the first 3D generative method that models the distribution of individual human geometry distributions. Specifically, we encode each human geometry distribution into a feature map, formulating 3D human generation as a two-stage process: first, the human geometry distribution (i.e., the feature map) is generated, and then, a high-fidelity human is synthesized by sampling from this distribution. Unlike the prior geometry distributions [52], which learn transformations from Gaussian noise to a selected shape, our key insight is to leverage the SMPL model distribution, which is inherently closer to human geometry. Instead of randomly sampling points from the SMPL model for training, we construct a mapping based on the proximity between points on the SMPL model and the human geometry surface, avoiding the inefficiencies of learning transformations between distant points. Furthermore, based on this mapping, we normalize both the SMPL and human distributions to reduce modeling complexity further. Given the distribution formulation, an auto-decoder encodes each geometry into a compact feature map, which, when combined with a human pose, represents a fine-grained human geometry capable of capturing realistic clothing details that naturally conform to body articulation.
@@
Our method synthesizes high-fidelity geometry directly, without relying on techniques like normal maps to enhance rendering results. We design two generative tasks to evaluate the effectiveness of our method. The first task generates random 3D human geometries conditioned on a given pose, while the second enables novel pose generation from a single-view normal image. Our quantitative results show that our method improves the geometry quality by $57\%$ ($42.9$ to $16.2$) and the visual appearance by $7\%$ ($17.4$ to $16.2$) compared to the state-of-the-art. Our contribution can be summarized as follows:
  • We propose the first 3D generative method that models the distribution of individual human geometry distributions, synthesizing high-fidelity avatars with realistic clothing details, including loose parts.
  • We introduce a novel framework for modeling human geometry distributions that samples from the SMPL model, significantly reducing modeling complexity and improving training efficiency.
  • We demonstrate the effectiveness of our geometry distribution in downstream tasks, achieving superior performance compared to existing generative methods.
@@ 2.1 General 3D Generation
3D generation methods employ various representations, such as meshes, voxels, point clouds, and implicit functions, each offering distinct advantages but also facing inherent limitations [52]. Here, we focus on compact representations that scale efficiently to large datasets.
Given one or more images, direct inference of pixel-aligned positions [48, 32, 43, 42] avoids intermediate representations, which is unsuitable for modeling intrinsic human shapes. A popular alternative is to synthesize tri-planes as intermediate representations [33, 19, 12], from which implicit functions or points [60, 9] can be sampled. However, tri-planes are limited by their resolution, restricting their ability to capture fine-grained details. GaussianCube [51] transfers Gaussians into a regular voxel grid via optimal transport, limiting its flexibility for intricate articulation. Gamba [31] and VectorSet [50] employ cross-attention mechanisms to inject 3D embeddings or image tokens, albeit constrained by memory consumption. Trellis [37] leverages sparse voxels but lacks mechanisms to incorporate human-specific priors. In contrast, 2D maps [25, 7, 46], offer a memory-efficient alternative that aligns naturally with human priors.
@@
Features extracted from images [26, 27, 2] provide rich view-dependent information, enabling high-fidelity human reconstruction. Some works incorporate additional cues such as voxel grids [10], human templates [59], or multi-view images [30] to mitigate the depth ambiguity of single-view image. For example, ICON [39] and ECON [40] predict the frontal and back normals from the SMPL template and reconstruct the human using these maps. Building on these reconstruction techniques, generative networks—such as GANs [38, 14] and diffusion models [29, 30, 54]—have been employed to synthesize pixel-aligned features or generate multi-view images [18] for human generation. However, these methods are inherently limited by their reliance on view-dependent features, sensitivity to multi-view inconsistencies, and often require post-processing to fuse information from different views [54], failing to capture the intrinsic 3D shape.
To learn intrinsic 3D representations, a popular approach [24, 53, 47, 35, 36] is to train a GAN model that synthesizes tri-planes and uses NeRF for rendering. However, these methods are constrained by rendering speed and resolution, often requiring super-resolution modules [3, 6], multi-part structures [45], or refinement stages [23, 58] to achieve high-fidelity results. Rather than employing tri-planes, alternative methods incorporate human templates, alleviating sparse spatial occupancy, but struggle with loose clothing [11] or still rely on super-resolution modules [13]. Explicit representations, such as primitive volumes [5], provide efficient rendering but face challenges in extracting detailed geometry. Mesh-based methods, using displacement [28] or layered surface volumes [44], are limited by fidelity and struggle with loose clothing. Point-based representations [1, 56, 55] offer flexibility for modeling thin structures but are constrained by point density. Furthermore, it can be difficult to produce accurate 3D geometry under indirect supervision from rendering. In contrast, gDNA [41] proposes learning implicit functions directly from 3D data, but synthesizing high-fidelity geometry remains challenging due to inherent representation limitations [52]. Our approach directly learns from 3D data while enabling “infinite” sampling, leading to high-fidelity human geometry synthesis. See Tab. 1.
@@
Tri-plane + NeRF [24, 53]
Template-based [11, 13]
Primitive Volumes [5]
Mesh-based [28, 44, 8]
Point-based [1, 56]
Implicit Function [41, 57]
@@ 3 Preliminaries
3.1 Flow Matching
Flow matching [20] is a variant of diffusion models. It constructs a flow that transforms samples from a source distribution $\mathbf{p}$ to a target distribution $\mathbf{q}$. Specifically, the training objective loss is defined as follow:
@@ 3.1 Flow Matching
3.2 Geometry Distributions
Geometry distributions model a surface $\mathcal{M}\subset\mathbb{R}^{3}$ as a probability distribution $\Phi_{\mathcal{M}}$, where any sample $\mathbf{x}_{1}\sim\Phi_{\mathcal{M}}$ corresponds to a point on the surface. To achieve this, Zhang et al. [52] adapt a diffusion model to learn the transformation from a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$ to the target distribution of surface points. A point $\mathbf{x}_{1}\in\mathcal{M}$ on the target surface can then be obtained by solving an ODE with an initial point $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@ 4.1 Overview
We model human geometry as a distribution based on the SMPL model and encode it into a compact feature map, which enables efficient generation within the dataset distribution. In this section, we first introduce our human geometry distribution formulation by identifying suitable source and target distributions. Next, we describe encoding this distribution into a feature map using an auto-decoder architecture, as depicted in Fig. 11 (a). Finally, we present generative tasks aimed at synthesizing the feature map, guided by the SMPL template, optionally incorporating additional conditioning inputs, as illustrated in Fig. 11 (b).
@@ 4.2 Human Geometry Distribution
Our insight is to replace the Gaussian distribution $\mathcal{N}(0,1)$ in the prior geometry distributions [52] with the SMPL template shape distribution $\Phi_{\mathcal{S}}$, aligning the source distribution closer to the target geometry distribution $\Phi_{\mathcal{T}}$. A naive approach can be formulated as:
@@
Training Pairs Construction.
Since the learned probability flow by Eq. 2 approximates a conditional optimal transport [21], one strategy for improving training efficiency is avoiding learning extraneous paths between distant points. To achieve this, we prioritize short flow by constructing training pairs set $\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ where $\mathbf{x}^{\prime}_{0}\sim\Phi_{\mathcal{S}}$ is geometrically close to $\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}$. Specifically, we first sample a set of points $\{\mathbf{x}_{0}\}_{\mathcal{S}}$ on the SMPL template and a set of points $\{\mathbf{x}_{1}\}_{\mathcal{T}}$ on the target geometry. Then, for each $\mathbf{x}_{1}\in\{\mathbf{x}_{1}\}_{\mathcal{T}}$, the corresponding $\mathbf{x}_{0}^{\prime}$ is obtained by:
$\mathbf{x}^{\prime}_{0}=\arg\min_{\mathbf{x}_{0}\in\{\mathbf{x}_{0}\}_{\mathcal{S}}}||\mathbf{x}_{1}-\mathbf{x}_{0}||_{2}.$
Notably, because multiple points $\mathbf{x}_{1}$ may share the same nearest SMPL points $\mathbf{x}_{0}^{\prime}$, especially in loose or wrinkled regions, directly learning a path from $\mathbf{x}_{0}^{\prime}$ to $\mathbf{x}_{1}$ leads to under-sampled geometry. Therefore, we add a perturbation drawn from $\mathcal{N}(0,\boldsymbol{\sigma})$ to $\mathbf{x}_{0}^{\prime}$, introducing randomness to enhance sample diversity. In this formulation, the source distribution becomes $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ and the target distribution is the human geometry. Sampling from the constructed training pairs, the flow matching optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma}),(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}}\left\|\cdot\right\|$ (3)
@@
Learning the transformation from the SMPL template to the human geometry of a dataset yields a spatial imbalance of sampling points, which arises from two factors. First, regions like the torso exhibit high point density because they remain in similar locations across different poses, while other regions are more dynamic and thus sparsely sampled. Second, the points are distributed on 2D subspaces within 3D space, leading to inherently sparse coverage. This imbalance leads to inefficient training as the model underfits low-density regions. To address this, we normalize both the source and target distribution by subtracting $\mathbf{x}_{0}^{\prime}$. The source distribution becomes a zero-centered Gaussian distribution $\mathcal{N}(0,1)$, with $\boldsymbol{\sigma}$ set to $1$, and the target distribution is modeled as a regularized dense displacement field $\Delta\mathbf{x}=\mathbf{x}_{1}-\mathbf{x}^{\prime}_{0}$. Note that this subtraction removes the positional information of $\mathbf{x}^{\prime}_{0}$. To retain this information while avoiding the imbalance impact, we reintroduce $\mathbf{x}_{0}^{\prime}$ as a conditional signal that scales the network hidden feature, indirectly influencing the feature representation. By building flow between regularized dense spaces, we reduce the modeling complexity and achieve improved training efficiency, as demonstrated in Sec 5.1.2. The optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime})-(\Delta\mathbf{x}-\mathbf{n})\right\|,\vspace{-0.5em}$ (4)
@@
Figure 3: Overview of our method. (a) We encode a geometry into a feature map, which is decompressed with a SMPL vertex map. The decompressed representation serves as a condition for our denoising network. (b) The human generation task is formulated as the conditional generation of feature maps, guided by the SMPL vertex map, optionally incorporating additional conditioning inputs.

### v3

@@ Abstract
Realistic human geometry generation is an important yet challenging task, requiring both the preservation of fine clothing details and the accurate modeling of clothing-body interactions. To tackle this challenge, we build upon Geometry distributions—a recently proposed representation that can model a single human geometry with high fidelity using a flow matching model. However, extending a single-geometry distribution to a dataset is non-trivial and inefficient for large-scale learning. To address this, we propose a new geometry distribution model by two key techniques: (1) encoding distributions as 2D feature maps rather than network parameters, and (2) using SMPL models as the domain instead of Gaussian and refining the associated flow velocity field. We then design a generative framework adopting a two-staged training paradigm analogous to state-of-the-art image and 3D generative models. In the first stage, we compress geometry distributions into a latent space using a diffusion flow model; the second stage trains another flow model on this latent space. We validate our approach on two key tasks: pose-conditioned random avatar generation and avatar-consistent novel pose synthesis. Experimental results demonstrate that our method outperforms existing state-of-the-art methods, achieving a $57\%$ improvement in geometry quality.
@@ 1 Introduction
3D human geometry generation is a critical yet challenging task. The human body exhibits high-frequency clothing details, which are inherently difficult to synthesize. Therefore, the core of this task is the design of a representation that can capture both the underlying structure and fine details, posing two key challenges: 1) encoding high-frequency details into a low-dimensional manifold without losing fidelity; 2) modeling the relationship between clothing wrinkles and poses to preserve realistic details.
Figure 1: Geometry distribution.
@@
Existing methods represent the human body in different ways. NeRFs Men et al. (2024); Zhang et al. (2022); Xiong et al. (2023); Jiang et al. (2023); Xu et al. (2023); Zheng et al. (2024) focus on the rendering results but neglect underlying geometry and are constrained by the rendering speed and resolution. Implicit functions Xu et al. (2022), such as signed distance functions, struggle to synthesize thin structures and tend to oversmooth results. Point clouds Zhang et al. (2024c; d) and volume-based representations trade off between memory efficiency and quality. Recently, geometry distributions Zhang et al. (2025) model single 3D shapes as distributions of points on their surfaces. Samples $\mathbf{X}$ from a Gaussian distribution $\mathcal{N}$ are transformed via a flow-based diffusion network $\mathcal{E}(\mathbf{X})$ into the target geometry, capturing fine-grained surface details (see Fig. 1). This formulation enables infinite point sampling, allowing for a high-fidelity representation of individual shapes. However, extending a single-geometry distribution to an entire dataset is non-trivial because of two reasons: 1) single-geometry formulations store the geometry in the parameters of a flow network, which leads to prohibitive memory consumption and limits their scalability for generative tasks, and 2) while learning the flow velocity fields from Gaussian distributions to a single shape is feasible, extending this to multiple shapes across a dataset becomes computationally expensive and inefficient.
Figure 2: (a) Denoising process of a human geometry distribution. (b) Random avatar generation for a given pose. (c) Novel pose generation of a given avatar. Results are rendered from point clouds.
To this end, we propose a novel human geometry distributions model by two key techniques:
  • We encode each human geometry distribution into a feature map rather than the network weights of a flow network directly, providing a generalized way to represent geometry distributions.
  • We adopt the SMPL Pavlakos et al. (2019) model distribution instead of a Gaussian and refine the associated flow velocity field for efficient learning. Specifically, we construct spatially efficient mappings from the velocity field and reformulate them into a regularized dense space.
@@
Based on these designs, we propose the first 3D generative method for geometry distributions, modeling the distribution of individual human geometry distributions. Our framework is formulated in two stages: first, we use a flow model to compress each human geometry distribution into a compact feature map, from which a high-fidelity human geometry can be sampled through a denoising process (Fig. 2 (a)), and then, we train another flow model to generate human geometry distributions (i.e., the feature maps). While the currently top-performing methods Xu et al. (2022); Zhang et al. (2024d; 2023b) obtain visually plausible geometric details by enhanced rendering techniques, our method synthesizes high-fidelity geometry directly, leading to large improvements. We design two generative tasks to evaluate the effectiveness of our method (Fig. 2 (b) and (c)). The first task generates random 3D human geometries conditioned on a given pose, while the second enables novel pose generation of a given avatar. Our quantitative results show that our method improves the geometry quality by $57\%$ ($42.9$ to $16.2$) and the visual appearance by $7\%$ ($17.4$ to $16.2$) compared to the state-of-the-art.
@@ 2.1 General 3D Generation
3D generation methods employ various representations, such as meshes, voxels, point clouds, and implicit functions, each offering distinct advantages but also facing inherent limitations Zhang et al. (2025). Here, we focus on compact representations that scale efficiently to large datasets.
Given one or more images, direct inference of pixel-aligned positions Yu et al. (2021a); Tang et al. (2024); Xu et al. (2024b; a) avoids intermediate representations, which is unsuitable for modeling intrinsic human shapes. A popular alternative is to synthesize tri-planes as intermediate representations Wang et al. (2023); Li et al. (2023); Hong et al. (2023), from which implicit functions or points Zou et al. (2024); Han et al. (2024) can be sampled. However, tri-planes are limited by their resolution, restricting their ability to capture fine-grained details. GaussianCube Zhang et al. (2024a) transfers Gaussians into a regular voxel grid via optimal transport, limiting its flexibility for intricate articulation. Gamba Shen et al. (2024) and VectorSet Zhang et al. (2023a) employ cross-attention mechanisms to inject 3D embeddings or image tokens, albeit constrained by memory consumption. Trellis Xiang et al. (2024) leverages sparse voxels but lacks mechanisms to incorporate human-specific priors. In contrast, 2D maps Rai et al. (2025); Elizarov et al. (2024); Yan et al. (2024), offer a memory-efficient alternative that aligns naturally with human priors.
@@
Features extracted from images Saito et al. (2019; 2020); Alldieck et al. (2022) provide rich view-dependent information, enabling high-fidelity human reconstruction. Some works incorporate additional cues such as voxel grids He et al. (2020), human templates Zheng et al. (2021), or multi-view images Shao et al. (2022) to mitigate the depth ambiguity of single-view image. For example, ICON Xiu et al. (2022) and ECON Xiu et al. (2023) predict the frontal and back normals from the SMPL template and reconstruct the human using these maps. Building on these reconstruction techniques, generative networks—such as GANs Xiong et al. (2023); Jiang et al. (2023) and diffusion models Sengupta et al. (2024); Shao et al. (2022); Zhang et al. (2024b)—have been employed to synthesize pixel-aligned features or generate multi-view images Kolotouros et al. (2024) for human generation. However, these methods are inherently limited by their reliance on view-dependent features, sensitivity to multi-view inconsistencies, and often require post-processing to fuse information from different views Zhang et al. (2024b), failing to capture the intrinsic 3D shape.
To learn intrinsic 3D representations, a popular approach Noguchi et al. (2022); Zhang et al. (2022); Yang et al. (2024); Wu et al. (2023; 2024) is to train a GAN model that synthesizes tri-planes and uses NeRF for rendering. However, these methods are constrained by rendering speed and resolution, often requiring super-resolution modules Bergman et al. (2022); Dong et al. (2023), multi-part structures Xu et al. (2023), or refinement stages Men et al. (2024); Zheng et al. (2024) to achieve high-fidelity results. Rather than employing tri-planes, alternative methods incorporate human templates, alleviating sparse spatial occupancy, but struggle with loose clothing Hong et al. (2022) or still rely on super-resolution modules Hu et al. (2024). Explicit representations, such as primitive volumes Chen et al. (2023), provide efficient rendering but face challenges in extracting detailed geometry. Mesh-based methods, using displacement Sanyal et al. (2024) or layered surface volumes Xu et al. (2024c), are limited by fidelity and struggle with loose clothing. Point-based representations Abdal et al. (2024); Zhang et al. (2024d; c) offer flexibility for modeling thin structures but are constrained by point density. Furthermore, it can be difficult to produce accurate 3D geometry under indirect supervision from rendering. In contrast, gDNA Xu et al. (2022) proposes learning implicit functions directly from 3D data, but synthesizing high-fidelity geometry remains challenging due to inherent representation limitations Zhang et al. (2025). Our approach directly learns from 3D data while enabling “infinite” sampling, leading to high-fidelity geometry synthesis. See Tab. 1 for comparison.
Table 1: Comparison of 3D human representation methods.
@@
Tri-plane + NeRF Noguchi et al. (2022); Zhang et al. (2022)
Template-based Hong et al. (2022); Hu et al. (2024)
Primitive Volumes Chen et al. (2023)
Mesh-based Sanyal et al. (2024); Xu et al. (2024c); Feng et al. (2023)
Point-based Abdal et al. (2024); Zhang et al. (2024d)
Implicit Function Xu et al. (2022); Zhang et al. (2023b)
@@ 3 Preliminaries
Flow Matching
Flow matching Lipman et al. (2023) is a variant of diffusion models. It constructs a flow that transforms samples from a source distribution $\mathbf{p}$ to a target distribution $\mathbf{q}$. Specifically, the training objective loss is defined as follow:
@@
Geometry Distributions
Geometry distributions model a surface $\mathcal{M}\subset\mathbb{R}^{3}$ as a probability distribution $\Phi_{\mathcal{M}}$, where any sample $\mathbf{x}_{1}\sim\Phi_{\mathcal{M}}$ corresponds to a point on the surface. To achieve this,  Zhang et al. (2025) adapt a diffusion model to learn the transformation from a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$ to the target distribution of surface points. A point $\mathbf{x}_{1}\in\mathcal{M}$ on the target surface can then be obtained by solving an ODE with an initial point $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@ 4.1 Overview
In this section, we first introduce our human geometry distribution formulation by identifying suitable source and target distributions (Sec. 4.2). Next, we describe encoding this distribution into a feature map using an auto-decoder architecture, as depicted in Fig. 4 (a) (Sec. 4.3). Finally, we present generative tasks aimed at synthesizing the feature map, guided by the SMPL template, optionally incorporating additional conditioning inputs, as illustrated in Fig. 4 (b) (Sec. 4.4).
@@ 4.2 Human Geometry Distribution
Our insight is to replace the Gaussian distribution $\mathcal{N}(0,1)$ in the prior geometry distributions Zhang et al. (2025) with the SMPL template shape distribution $\Phi_{\mathcal{S}}$, aligning the source distribution closer to the target geometry distribution $\Phi_{\mathcal{T}}$. A naive approach can be formulated as:
@@
Training Pair Construction.
Since the learned probability flow by Eq. 2 approximates a conditional optimal transport Lipman et al. (2024), one strategy for improving training efficiency is avoiding learning extraneous paths between distant points. To achieve this, we prioritize short flow by constructing training pairs set $\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ where $\mathbf{x}^{\prime}_{0}\sim\Phi_{\mathcal{S}}$ is geometrically close to $\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}$. Specifically, we first sample a set of points $\{\mathbf{x}_{0}\}_{\mathcal{S}}$ on the SMPL template and a set of points $\{\mathbf{x}_{1}\}_{\mathcal{T}}$ on the target geometry. Then, for each $\mathbf{x}_{1}\in\{\mathbf{x}_{1}\}_{\mathcal{T}}$, the corresponding $\mathbf{x}_{0}^{\prime}$ is obtained by:
[…]
Notably, because multiple points $\mathbf{x}_{1}$ may share the same nearest SMPL points $\mathbf{x}_{0}^{\prime}$, especially in loose or wrinkled regions, directly learning a path from $\mathbf{x}_{0}^{\prime}$ to $\mathbf{x}_{1}$ leads to under-sampled geometry (similar to the artifacts shown in the second row of Fig. 7). Therefore, we add a perturbation drawn from $\mathcal{N}(0,\bm{\sigma})$ to $\mathbf{x}_{0}^{\prime}$, introducing randomness to enhance sample diversity. In this formulation, the source distribution becomes $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ and the target distribution is the human geometry. Sampling from the constructed training pairs, the flow matching optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma}),(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}}\left\|\cdot\right\|$ (3)
@@
Figure 3: Aggregated samples $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})$.
When training the transformation from the SMPL template to the human geometry of a dataset (Eq. 3), the network receives spatially imbalanced supervision due to unevenly sampling points, which arises from two factors. First, points lie only on the surfaces, making them extremely sparse and unevenly distributed relative to the full 3D space. Second, variations in human pose and body shape further exacerbate this spatial imbalance, as some regions are more dynamic and thus sampled less consistently across geometries. Fig. 3 visualizes this effect by aggregating sampled points from multiple training iterations, showing that some spatial regions have high point density while others remain sparsely sampled. This spatial imbalance leads to inefficient training, as the model pays more attention to high-density regions while neglecting low-density regions. To address this, we normalize both the source and target distribution by subtracting $\mathbf{x}_{0}^{\prime}$. The source distribution becomes a zero-centered Gaussian distribution $\mathcal{N}(0,1)$, with $\bm{\sigma}$ set to $1$, and the target distribution is modeled as a regularized dense displacement field $\Delta\mathbf{x}=\mathbf{x}_{1}-\mathbf{x}^{\prime}_{0}$. Note that this subtraction removes the positional information of $\mathbf{x}^{\prime}_{0}$. To retain this information without suffering from imbalanced sampling, we reintroduce $\mathbf{x}_{0}^{\prime}$ as a conditional signal that scales the network hidden feature, indirectly influencing the feature representation. By building flow between regularized dense spaces, we reduce the modeling complexity and achieve improved training efficiency, as demonstrated in Sec 5.2. The optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime})-(\Delta\mathbf{x}-\mathbf{n})\right\|,\vskip-5.0pt$ (4)
@@
Figure 4: Overview of our method. (a) We encode a geometry into a feature map, which is decompressed with a SMPL vertex map. The decompressed feature serves as a condition for our denoising network. (b) The human generation task is formulated as the conditional generation of feature maps, guided by the SMPL vertex map, optionally incorporating additional conditioning inputs.